# Avoiding matrix exponentials for large transition rate matrices

### 09 Dec 2023 v1

@@ A.4.3 Parameters in the adaptive scheme
In the case of the birth-death process, the initialization guarantees the sampler starts in a region of high posterior probability. Consequently, we terminate the updating procedure after completing 10 cycles, each consisting of 100 samples, where the acceptance rate —defined as the ratio of accepted proposals—falls within the range of $0.2$ to $0.5$. Note that the counting of cycles is independent of the updating of $\mathbf{Z}$ at each 300 samples, thus the sampler has to “survive” at least 3 changes in $\mathbf{Z}$ by keeping a reasonable acceptance rate before settling.
@@ Appendix B Consistency Results
In this Supplemental Information section we present samples from the posterior for all kinetic parameter values treated in Sec. 2.2 but not plotted in the main text. The posterior samples histograms for the birth-death process (Sec. 3.1.1) are found in Fig. 6, the ones for the two state birth-death (Sec. 3.1.2) are found in Fig. 7, while the autoregulating gene example (Sec. 3.1.3) is found in Fig. 8.
To ensure that one can obtain sufficient samples to fully characterize the posterior, $p(\theta|\bar{s})$ in feasible time we obtain $250000$ samples of the kinetic parameters, $\theta$, for each dataset analyzed. Density histograms are then constructed from these samples. Each histogram focuses on individual elements of $\theta$, e.g., for the birth-death process we have $\theta=\{\beta,\gamma\}$ and the density histogram for the sampled values of $\beta$ in an approximation of the marginalized posterior – $p(\beta|\bar{s})=\int\mathrm{d}\gamma\ p(\theta=\{\beta,\gamma\}|\bar{s})$. These histograms are juxtaposed with the sample average (approximating the marginal posterior expected value) and the 50 % and 95% credible intervals.
Figure 6: Density histogram of samples from the posterior distribution for the birth-death process (Sec. 3.1.1). Different panels represent posterior obtained with datasets of varying kinetic parameter.
Figure 7: Density histogram of samples from the posterior distribution for the two state birth-death process (Sec. 3.1.2). Different panels represent the posterior obtained with datasets of varying kinetic parameters.
Figure 8: Density histogram of samples from the posterior distribution for the autoregulating gene process (Sec. 3.1.3). Different panels represent the posterior obtained with datasets of varying kinetic parameters.
d7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 06 Feb 2024 v3

@@ A.4.3 Parameters in the adaptive scheme
In the case of the birth-death process, the initialization guarantees the sampler starts in a region of high posterior probability. Consequently, we terminate the updating procedure after completing 10 cycles, each consisting of 100 samples, where the acceptance rate (defined as the ratio of accepted proposals) falls within the range of $0.2$ to $0.5$. Note that the counting of cycles is independent of the updating of $\mathbf{Z}$ at each 300 samples, thus the sampler has to “survive” at least 3 changes in $\mathbf{Z}$ by keeping a reasonable acceptance rate before settling.
@@ Appendix B Consistency Results
In this Supplemental Information section we present samples from the posterior for all kinetic parameter values treated in Sec. 2.2 but not plotted in the main text. The posterior sample histograms for the birth-death process (Sec. 3.1.1) are found in Fig. A1, the ones for the two state birth-death (Sec. 3.1.2) are found in Fig. LABEL:fig:SI2S, while the autoregulating gene example (Sec. 3.1.3) is found in Fig. A3.
To ensure that one can obtain sufficient samples to fully characterize the posterior, $p(\theta|\bar{s})$ in feasible time we obtain $250000$ samples of the kinetic parameters, $\theta$, for each dataset analyzed. In the J-MJP, we remove all other variables keeping only the samples of the kinetic parameters ( which is equivalent to marginalizing these latent variables). The density histograms are then constructed from these samples. Each histogram focuses on individual elements of $\theta$, e.g., for the birth-death process we have $\theta=\{\beta,\gamma\}$ and the density histogram for the sampled values of $\beta$ in an approximation of the marginalized posterior, $p(\beta|\bar{s})=\int\mathrm{d}\gamma\ p(\theta=\{\beta,\gamma\}|\bar{s})$. These histograms are juxtaposed with the sample average (approximating the marginal posterior expected value) and the 50 % and 95% credible intervals. See Fig. 4 in the main text for all conventions.
Figure A1: Density histogram of samples from the posterior distribution for the birth-death process (Sec. 3.1.1). Different panels represent posterior obtained with datasets of varying kinetic parameter. See Fig. 4 in the main text for conventions.
Figure A2: Density histogram of samples from the posterior distribution for the two state birth-death process (Sec. 3.1.2). Different panels represent the posterior obtained with datasets of varying kinetic parameters. See Fig. 4 in the main text for conventions.
Figure A3: Density histogram of samples from the posterior distribution for the autoregulating gene process (Sec. 3.1.3). Different panels represent the posterior obtained with datasets of varying kinetic parameters. See Fig. 4 in the main text for conventions.